Problem #7: A certain system can experience three different types of defects. Let Ai (1 1,2,3) denote the event that the system has a defect of type i. Suppose that P(41) = 0.3 l, P(d2) = 0.29, Padg) = 0.35, P(A2 U A3) = 0.59. P(41 n.42 n A3) = 0.03 (a) Find the probability that the system has exactly 2 of the 3 types of defects. (b) Find the probability that the system has a type 1 defect given that it does not have a type 2 or type 3 defect. Problem #7(a): Problem #7(b):
Answer #1

A)

Probability for exactly 2 of 3 type of defects. = P ( A1 \cap A2 ) + P ( A1 \cap A3 ) + P ( A3 \cap A2 ) - 2 P ( A1 \cap A2 \cap A3 )

P ( A1 \cap A2 ) = P(A1) + P(A2) - P ( A1 U A2 ) = 0.07

P ( A2 \cap A3 ) = P(A2) + P(A3) - P ( A2 U A3 ) = 0.05

P ( A1 \cap A3 ) = P(A1) + P(A1) - P ( A1 U A3 ) = 0.09

Probability for exactly 2 of 3 type of defects. = 0.07 + 0.05 + 0.09 - 0.03

= 0.18

B)

Type 1 effect but not type 2 and 3 = P ( A1  \capA2 ) + P ( A1  \capA3 ) / P(A1)

= ( 0.09 + 0.07 )/ 0.31

= 0.5161